The density of titanium metal is 4.51 g/cm3 at 25 ∘C. What mass of titanium displaces 64.2 mL of water at 25 ∘C?


1
Expert's answer
2020-09-08T05:57:23-0400

1 ml of water is 1 cm3

"V_{displased} = 64.2 cm^3"

Density "ρ = \frac{m}{V}"

"m = ρ\times V"

"m = 4.51\frac{g}{cm^3} \times 64.2 cm^3= 289.54 g"

Answer: 289.54 g